OpenStudy (anonymous):

@satellite73 last one. :)

OpenStudy (anonymous):

whew it is past my curfew

OpenStudy (anonymous):

Use the following table to answer questions 1 and 2. The two-way table shows the number of students in a middle school who have 0, 1, 2, or 3 siblings. Siblings Grade 7 Grade 8 Total 0 10 7 17 1 15 14 29 2 8 18 26 3 7 6 13 Total 40 45 85 How many more 7th graders than 8th graders have 1 or no siblings? A. 1 B. 3 C. 4 <----- D. 12 @satellite73 @mathslover @mathmale

OpenStudy (anonymous):

seventh grades with 0 or 1 sibling: 10+15=25

OpenStudy (anonymous):

eigths graders with 0 or 1 sibling : 7 + 14 = 21

OpenStudy (anonymous):

\(25-21=4\) you win again